Before Eid-UL-Fitr, the price of juice was Rs.100 a box in Karachi, and Hayat Khan was willing to buy 10 juice boxes. After the Eid-UL-Fitr holidays, the price has gone up to Rs.120 a box, and Khan is now willing to buy 8 juice boxes. Is khan’s demand for juice boxes, elastic or inelastic? What is Khan’s elasticity of demand? Draw Khan’s demand curve, label everything clearly.


1
Expert's answer
2021-12-05T18:57:46-0500

The elasticity of demand is used to determine how responsive is the quantity demanded of a good when there is a change in the price of that good. 

 

In the given case, the elasticity of demand for juice is:

The elasticity of demand for juice

"=\frac{Change \space in\space quantity\space demanded}{Change\space in\space price} \times \frac{Initial Price}{Initial quantity}\\=\frac{ −2}{20} \times \frac{100}{10}\\=1"

Yes, the demand for juice boxes is elastic. 

 

The elasticity of demand is 1. It is unitary elastic.